OpenStudy (anonymous):

Solve for the missing side in the following right triangle: hypotenuse = 17, leg = 8, leg = ?.

OpenStudy (anonymous):

a^2 +c^2 = c^2 8^2+x^2 = 17^2 solve

OpenStudy (mathstudent55):

Have you learned the Pythagorean theorem? You need it here.

OpenStudy (anonymous):

a^2 +b^2 = c^2 typo

OpenStudy (anonymous):

15? Yeah I have learned it but I am not very good at it.

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):